# Sparse but Wrong: Incorrect L0 Leads to Incorrect Features in Sparse Autoencoders

### v2

@@ 4.2 Can L0 be both too low and too high simultaneously?
In Figure 10 (right), we plot decoder histogram projection plots for BatchTopK SAEs on Gemma-2-2b layer 12 with L0 10, 200, 750, and 2000. As we expected from Figure 6, when L0 is very low (10) or very high (2000), we see a wide gaussian around 0, indicating that decoder latents are mixing correlated features together. At L0=200, we see a much more narrow distribution around 0, as we expect when near the correct L0. However, at L0=750, we see an interesting phenomenon, where there is an even narrower distribution than at L0=200, but also a large hump starting at projection above 10 (more visible in the log plot).
We suspect this indicates at L0=750, some latents become more monosemantic while other latents mix underlying features becoming less monosemantic. This likely means that the L0 is too high for some latents while simultaneously being too low for other latents. There is no reason why every latent has the same optimal L0 threshold, so there is likely a range of L0s where some latents are firing more than they ideally should while other latents are firing less than they ideally should. We also suspect this is part of why JumpReLU SAEs seem to perform much better at high L0, since JumpReLU SAEs can adjust firing threshold per-latent while BatchTopK SAEs cannot.
@@ 5 Related work
Chanin et al. (2025) explores feature hedging, showing SAEs mix correlated features into latents if the SAE is too narrow. We consider our work a version of feature hedging due to low L0. Till (2024) shows SAEs may increase sparsity by inventing features. Chanin et al. (2024) discuss the problem of feature absorption, where SAEs can improve their sparsity score by mixing hierarchical features together. Engels et al. (2024) investigates SAE errors and finds that SAE error may be pathological and non-linear. Engels et al. (2025) find that not all underlying LLM features themselves are linear, demonstrating circular embeddings of some concepts. Wu et al. (2025) and Kantamneni et al. (2025) both investigate empirical SAE performance, finding SAEs underperform relative to supervised baselines, but do not offer theoretical explanations as to why SAEs underperform.
@@ 6 Discussion
While most practitioners of SAEs understand that having too high L0 is problematic, our work shows that having too low of L0 is perhaps even worse. Our work has several important implications for the field. First, the L0 used by most SAEs is lower than it ideally should be, as a cursory search of open source SAEs on Neuronpedia (Lin, 2023) shows L0 less than 100 is very common even for SAEs trained on large models (see Appendix A.7). We further show that the sparsity–reconstruction tradeoff, as commonly discussed by most SAE papers (Cunningham et al., 2024; Gao et al., 2024; Rajamanoharan et al., 2024), is misleading: when L0 is too low, an SAE with a correct dictionary achieves worse reconstruction than an incorrect SAE that mixes correlated features.
We presented a metric based on the correlation between the SAE decoder and input activations, $s_{n}^{\text{\text{dec}}}$, that can give us hints about the correct L0 for a given SAE. However, we do not view this as a perfect guide. As we saw in our results, while low L0 SAEs consistently have very high $s_{n}^{\text{\text{dec}}}$, the metric is not always high at high L0 (although it usually is). Still, we feel that this metric is a useful guide to avoid L0 that is clearly too low, and we hope this investigation into correlation-based SAE quality metrics can be built on further in future work.
While our metric currently requires training a sweep over L0 to optimize, we are hopeful that it may be possible to optimize this metric automatically during training (steps towards this are discussed in Appendix A.6). We further encourage anyone implementing ideas from this work to also experiment with our alternative metric, decoder pairwise cosine similarity, discussed in Appendix A.5.
@@ Acknowledgments
David Chanin was supported thanks to EPSRC EP/S021566/1 and the Machine Learning Alignment and Theory Scholars (MATS) program. We are grateful to Henning Bartsch for feedback during the project.
@@ A.1 SAE training architecture definitions
In this work we focus on JumpReLU (Conerly et al., 2025; Rajamanoharan et al., 2024) and BatchTopK (Bussmann et al., 2024) SAEs. For BatchTopK SAEs, there is no sparsity penalty as sparsity is enforced by the BatchTopK function. The auxiliary loss $\mathcal{L}_{P}$ for BatchTopK is as follows, where $e$ is the SAE training error residual, and $\hat{e}$ is a reconstruction using the top $k_{\text{aux}}$ dead latents (meaning the latents have not fired in more than $n_{\text{dead}}$ steps during training).
@@
We follow the JumpReLU training setup from Conerly et al. (2025), which involves both a sparsity loss $\mathcal{L}_{s}$ and a pre-activation loss for reviving dead latents, $\mathcal{L}_{p}$. $\mathcal{L}_{s}$ is defined as below, where $c$ is a scalar scaling factor:

### v3

@@ 4.2 Can L0 be both too low and too high simultaneously?
In Figure 9 (right), we plot decoder projection histogram plots for BatchTopK SAEs on Gemma-2-2b layer 12 with L0 10, 200, 750, and 2000. These plots are created by projecting training inputs on the SAE decoder, creating a histogram of how strongly each latents projects onto the input. We expect that the more SAE latents are mixing positive and negative components of underlying features, the more strongly they should project both positively and negatively on arbitrary training inputs. This should look like a narrow gaussian around 0 when there is little mixing, and a wider gaussian the more mixing there is. This is also the intuition behind the alternative metric discussed in Appendix A.9, and the theory behind this is formalized further in Appendix A.10.
As expected, when L0 is very low (10) or very high (2000), we see a wide gaussian around 0, indicating that decoder latents are mixing correlated features together. At L0=200, we see a much more narrow distribution around 0, as we expect when near the correct L0. However, at L0=750, we see an interesting phenomenon, where there is an even narrower distribution than at L0=200, but also a large hump starting at projection above 10 (more visible in the log plot).
We suspect this indicates at L0=750, some latents become more monosemantic while other latents mix underlying features becoming less monosemantic. This likely means that the L0 is too high for some latents while simultaneously being too low for other latents. There is no reason why every latent has the same firing threshold, so there is likely a range of L0s where some latents are firing more than they ideally should while other latents are firing less than they ideally should. We also suspect this is part of why JumpReLU SAEs seem to perform much better at high L0, since JumpReLU SAEs can adjust firing threshold per-latent while BatchTopK SAEs cannot.
@@ 5 Related work
Limitations of SAEs
Early work on SAEs for interpretability highlight the problem of feature splitting (bricken2023towards; templeton2024scaling), where a seemingly interpretable general feature splits into more specific features at narrower SAE widths. chanin2025feature explores feature hedging, showing SAEs mix correlated features into latents if the SAE is too narrow. We consider our work a version of feature hedging due to low L0. till2024sparse shows SAEs may increase sparsity by inventing features. chanin2024absorption discuss the problem of feature absorption, where SAEs can improve their sparsity score by mixing hierarchical features together. engels2024decomposing investigates SAE errors and finds that SAE error may be pathological and non-linear. engels2025not find that not all underlying LLM features themselves are linear, demonstrating circular embeddings of some concepts. wu2025axbench and kantamneni2025sparse both investigate empirical SAE performance, finding SAEs underperform relative to supervised baselines, but do not offer theoretical explanations as to why SAEs underperform.
Picking SAE hyperparameters
Related to our work is Minimum Description Lengths (MDL) SAEs (ayonrinde2024interpretability), which attempt to find reasonable choices for SAE width and L0 based on information theory. However, MDL SAEs assume that there is no inherently “correct” decomposition for LLM activations and no “correct” L0, and therefore does not attempt to find the underlying true features. Our work takes the opposite approach, starting from simple toy models with linear features and showing that if L0 is not set correctly the SAE decoder becomes corrupted.
Another SAE architecture which attempts to pick L0 heuristically is Approximate Feature Activation (AFA) SAEs (lee2025evaluating). AFA SAEs selects L0 adaptively at each input by assuming underlying true features are maximally orthogonal and selecting features until the feature norm is close to the input norm. While the L0 is not set directly in AFA SAEs, there is an extra loss hyperparameter that may modulate the resulting L0.
@@ 6 Discussion
While most practitioners of SAEs understand that having too high L0 is problematic, our work shows that having too low of L0 is perhaps even worse. Our work has several important implications for the field. First, the L0 used by most SAEs is lower than it ideally should be, as a cursory search of open source SAEs on Neuronpedia (neuronpedia) shows L0 less than 100 is very common even for SAEs trained on large models (see Appendix A.13). We further show that the sparsity–reconstruction tradeoff, as commonly discussed by most SAE papers (huben2024sparse; gao2024scaling; rajamanoharan2024jumping), is misleading: when L0 is too low, an SAE with a correct dictionary achieves worse reconstruction than an incorrect SAE that mixes correlated features.
We presented a metric based on the correlation between the SAE decoder and input activations, $c_{\text{dec}}$, that can give us hints about the correct L0 for a given SAE. However, we do not view this as a perfect guide. As we saw in our results, while low L0 SAEs consistently have very high $c_{\text{dec}}$, the metric can sometime remain nearly flat for a wide range of L0. Still, we feel that this metric is a useful guide to avoid L0 that is clearly too low, and we hope this investigation into correlation-based SAE quality metrics can be built on further in future work. We are particularly excited about the possibility that we can learn more about the underlying correlational structure between underlying features by studying correlations in the SAE decoder.
While our metric currently requires training a sweep over L0 to optimize, we are hopeful that it may be possible to optimize this metric automatically during training (steps towards this are discussed in Appendix A.11). Improving this further is left to future work.
@@ Acknowledgments
David Chanin was supported thanks to EPSRC EP/S021566/1 and the Machine Learning Alignment and Theory Scholars (MATS) program. We are grateful to Henning Bartsch and Lovkush Agarwal for feedback during the project.
@@ A.1 SAE training architecture definitions
In this work we focus on JumpReLU (conerly2025jumprelu; rajamanoharan2024jumping) and BatchTopK (bussmann2024batchtopk) SAEs. For BatchTopK SAEs, there is no sparsity penalty as sparsity is enforced by the BatchTopK function. The auxiliary loss $\mathcal{L}_{P}$ for BatchTopK is as follows, where $e$ is the SAE training error residual, and $\hat{e}$ is a reconstruction using the top $k_{\text{aux}}$ dead latents (meaning the latents have not fired in more than $n_{\text{dead}}$ steps during training).
@@
We follow the JumpReLU training setup from conerly2025jumprelu, which involves both a sparsity loss $\mathcal{L}_{s}$ and a pre-activation loss for reviving dead latents, $\mathcal{L}_{p}$. $\mathcal{L}_{s}$ is defined as below, where $c$ is a scalar scaling factor: